Dillon Labs has asked its financial manager to measure the cost of each specific type of capital as well as the weighted average cost of capital. The weighted average cost is to be measured by using the following​ weights: 40​% ​long-term debt, 25​% preferred​ stock, and 35​% common stock equity​ (retained earnings, new common​ stock, or​ both). The​ firm's tax rate is 25​%. Debt The firm can sell for ​$1010 a 19​-year, ​$1 comma 000​-par-value bond paying annual interest at a 7.00​% coupon rate. A flotation cost of 2.5​% of the par value is required. Preferred stock  7.00​% ​(annual dividend) preferred stock having a par value of ​$100 can be sold for ​$92. An additional fee of ​$3 per share must be paid to the underwriters. Common stock  The​ firm's common stock is currently selling for ​$70 per share. The stock has paid a dividend that has gradually increased for many​ years, rising from ​$2.75 ten years ago to the ​$4.70 dividend​ payment, Upper D 0​, that the company just recently made. If the company wants to issue new new common​ stock, it will sell them ​$1.50 below the current market price to attract​ investors, and the company will pay ​$3.00 per share in flotation costs.   a.  Calculate the​ after-tax cost of debt. b.  Calculate the cost of preferred stock. c.  Calculate the cost of common stock​ (both retained earnings and new common​ stock). d.  Calculate the WACC for Dillon Labs.
Answer #1
Annual average growth rate=((last value/First value)^(1/Time between 1st and last value)-1)*100
Annual Growth rate=((4.7/2.75)^(1/10)-1)*100
Annual Growth rate% = 5.51
Weight of equity = E/A
Weight of equity =
W(E)=0.35
Weight of debt = D/A
Weight of debt = 0.4
W(D)=0.4
Weight of preferred equity =1-D/A-E/A
Weight of preferred equity = =1-0.4 - 0.35
W(PE)=0.25
Cost of equity
As per DDM
Price-flotation cost = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
65.5-4.5 = 4.7 * (1+0.0551) / (Cost of equity - 0.0551)
c. Cost of equity% = 13.08
Cost of debt
                                         K = N
Bond Price -flotation cost =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                                          k=1
                                         K =19
1010-25 =∑ [(7*1000/100)/(1 + YTM/100)^k]     +   1000/(1 + YTM/100)^19
                                          k=1
YTM = 7.1467303822
a. After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 7.1467303822*(1-0.25)
= 5.36
b, cost of preferred equity
cost of preferred equity = Preferred dividend/price-flotation cost*100
cost of preferred equity = 7/(89-3)*100
=7.87
d. WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=5.36*0.4+13.08*0.35+7.87*0.25
WACC =8.69%